A positive integer is picked randomly from 51 to 60, inclusive.
what is the probability that it is divisible by both 2 and 3?

Answer by stanbon(75887)

A positive integer is picked randomly from 51 to 60, inclusive.
what is the probability that it is divisible by both 2 and 3?
-----
Ways to succeed: 54 ; 60
-----
P(divis by 2 and 3) = 2/10 = 1/5
======================================
Cheers,
Stan H.
===============